OpenStudy (anonymous):

The price of a suit decreased from $280 to $182 during a clearance sale. What is the percent of decrease in the price? Round to the nearest percent, if necessary. A. 98% B. 65% C. 54% D. 35%

OpenStudy (anonymous):

Subtract $182 from $280 to find the difference. $280-$182=$98 Write a proportion.\[\frac{ 98 }{ 280 }=\frac{ x }{ 100 }\]Solve for x Cross multiply 98 times 100: 98*100=9800 Divide by the other number, 280. 9800/280 = 35